# Constraints to Lorentz violation and ultrahigh-energy electrons in D-foamy space-times

### v1

@@ 1 Introduction and summary
The discovery of PeV photons reported by the Large High Altitude Air Shower Observatory (LHAASO) from PeVatrons LHAASO:2021gok ; Li:2021tcw ; LHAASO:2023uhj including the ancient Crab Nebula LHAASO:2021cbz , kicks off a new era in astronomy, providing a unique opportunity to probe the potential breakdown of Lorentz invariance (LI) He:2022gyk ; AlvesBatista:2023wqm arising from quantum gravity (QG) analysis. High-energy violation of this symmetry (i.e., Lorentz violation (LV)) leads to particle-energy ($E$) dependent physics, that is suppressed by the gravitational scale $M\sim M_{P}\approx 10^{19}$ GeV where $M_{P}$ is the Planck mass. As a result, for extreme energies reached in astrophysical processes, its consequences could be identified in the current observations. The question arises, then, what information we can get from LHAASO experiments of ultrahigh-energy (UHE) $\gamma$-rays and particularly, from $\sim 1.1$ PeV Crab photon. Recent studies He:2022jdl , carrying further a previous analysis Li:2022ugz by the present authors, pointed out that the (indirect) observations of UHE electrons assumed to produce the PeV emission of the Crab severely limit minimally suppressed LV dispersion corrections, i.e., of ${\cal O}(E/M)$, to both electrons and photons.
Such bounds were placed via considerations of the vacuum Cherenkov process, $e\rightarrow e\gamma$, which is forbidden in LI quantum electrodynamics (QED), since the speeds of fast moving charges are always less than that of photon propagation. With LV it can be Coleman:1997xq an important energy-loss channel for the most high-energetic particle. As often adopted in the literature to parameterize LV (see He:2022gyk ), also in the same spirit as He:2022jdl ; Li:2022ugz , a priori most phenomenological studies converge in assuming a generic, nonstandard dispersion relation:
@@
where $(E_{I},{\bf p}_{I})$ is the four-momentum associated with a given particle $I$ of (rest) mass $m_{I}$, and the modulus of the spatial momentum is sometimes denoted, below, by $p\coloneqq\lvert{\bf p}\rvert$. [N.B.: the notation follows prior studies of He:2022jdl ; Li:2022ugz .] $\ell_{I,n}\coloneqq s_{I,n}/(E_{\textrm{LV}I,n})^{n}\equiv-\xi_{I,n}/M^{n}$ defines the scale of the $n$th-order LV physics beneath, [^3] Above, the dimensionless LV parameters $\xi_{I,n}$ is defined to simplify some of the expressions below, albeit they may be absorbed into redefinition of the mass $M$, and, then give the (positive) LV energy scale, $E_{\textrm{LV}I,n}$ by keeping only the LV sign factor $s_{I,n}(=\pm 1)$ explicitly. while $\delta^{(I\,=\,e)}$ parameterize the strength of ${\cal O}(p^{2})$ electron LV that has been constrained to be extremely small, $\lvert\delta\rvert<10^{-20}$ Li:2022ugz , implying that for astrophysically relevant processes, ${\cal O}(1/M^{n})$ terms dominant.
For example, in a linear LV model ($n=1$, in which case this suffix $n$ is omitted) often considered, modified dispersion (1) induces an energy-dependent velocity of propagation in vacuo for both the electromagnetic waves, $\Delta v_{\gamma}\coloneqq\partial E_{\gamma}/\partial p_{\gamma}-1\simeq-\,(\ell E)_{\gamma}$ and electron waves, $\Delta v_{e}\simeq-[m_{e}^{2}/(2E_{e}^{2})+(\ell E)_{e}]$, assuming vanishing $\delta$, with $s_{\gamma/e}=+1$ ($-1$) for subluminal (superluminal) propagation relative to low-energy photons that are not supposed to be affected significantly by the LV effects, and that will travel at the conventional speed of light $c$ (already set to be unity). [^4] Such nontrivial propagation properties can be found from concrete models. For example within effective field theory approach to modified QED model with operators of dimension-5 are compatible with the case of exhibiting birefringence (i.e., having Myers:2003fd helicity-dependent $\ell_{\gamma/e}$) while (Liouville Ellis:1992eh ) string inspired foamy space-time ‘medium’ Mavromatos:2002re based on D-brane recoil Ellis:1999uh dictates a subluminal refraction for photons (and not electrons, c.f. $\ell_{e}=0$ at tree-level), see section 3. Aside from such propagation effects in vacuo there is the possibility of peculiar novel (threshold) reactions beyond the standard model, like $e\rightarrow e\gamma$, provided the energy–momentum conservation relation is kept intact, an assumption usually adopted in a purely phenomenological formulation of LV, e.g., (1), owing to a lack of any particular prior knowledge.
@@
Quite interestingly, the absence of hard Cherenkov threshold severely constrains subluminal photon dispersions with linear LV modification, an effect compatible with a previous hint of light-speed variation in gamma-ray burst (GRB) photon flight time delays Shao:2009bv , that require $E_{\textrm{LV}\gamma}\sim 10^{17}$ GeV, leading to a saturation of the latter. This motivates us to reconsider the bound from (3). Indeed, only after one knows how rapidly the drops in electron energy increase, can any kinematic constraint from threshold considerations be trusted. So far this issue has only been treated in specific models of LV Anselmi:2011ae ; Klinkhamer:2008ky ; Rubtsov:2012kb , while for general cases like that of (1) the reaction rate has not been clarified. It is this what we plan to slightly touch upon in the first part of this article. If above threshold, it is high, as we shall see, for setting restrictions to this type of LV it then suffices to simply perform threshold analyses as in refs. He:2022jdl ; Li:2022ugz , which will certainly give accurate results.
Nonetheless, we need to emphasize here that at this kinematical level, the allowance of Cherenkov effects in a vacuum necessitates both i) the introduction of LV-modified dispersions, as parameterized by (1) and, ii) the intactness of energy–momentum conservations which, within some scenarios of QG, require re-examination. A possible failure of either of these two postulates is likely to preclude again any possibilities of ‘electron decay’ via such an anomalous channel, as in the standard model (but for totally different reasons), despite the existence of LV in the electron and/or photon sector(s). This may indeed characterize the situations encountered in certain (solitonic-)string/brane models of space-time foam to which we pay particular attention later on in this work.
@@ 2.1 A note on the calculation of rates in Lorentz violation contexts
We consider it as instructive to remind the reader that employing the microscopic language of rates, LV effect in a complete treatment is threefold. First, of particular relevance for us, deviations of phase space of particle kinematics. Second, modifications of wave functionals of external legs and the third, changes in the vertices and propagators. Scattering matrices may then get modified for the latter two types of corrections (compared to the standard LI case). However, considering the modified form of the matrix element ${\cal M}$ would require the precise knowledge of the LV dynamics, for example (in the case of field-theoretical models) the novel Feynman (and, spin sum) rules, due to extra terms present in the corresponding Lagrangian, which is obviously absent in our phenomenological setup (1). Of course, one can plausibly assume at least in the low energy limit, the amplitude squared (summed and averaged over spins and/or polarizations):
¯$\overline{\sum}\lvert{\cal M}^{\textrm{LV}}(\{p\})\rvert^{2}=\overline{\lvert{\cal M}(\{p\})\rvert^{2}}\,\bigl{[}1+{\cal O}(E/M)\bigr{]},$ (4)
@@
On the other note, the aim of this section is restricted to a matter of illustration of the calculation of Cherenkov rate, in order to justify the estimates of bounds from kinematical thresholds, supplementing recently published studies He:2022jdl ; Li:2022ugz . It is for this purpose as well we base our analysis here in the frame which is the same as the one thereof, neglecting possible effects in (4). Nonetheless, as mentioned above, such reaction rates have been calculated in many earlier works, to which we refer the reader for details, and, where one considers in general a concrete theory so that the dynamical effects (c.f. (4)) can be treated properly. For example, modifications of phase-space integrals and to the functional form of external states were studied in Jacobson:2005bg in the context of QED with dimension-5 field operators of which the dispersion induced is of $n=1$ type Myers:2003fd thus bears a similarity with the one considered here. Later, spin sums over external states for models with operators of higher dimensions were noted in Anselmi:2011ae , which built on earlier calculations of Klinkhamer:2008ky  [^6] Albeit severely constrained via the absence of detectable birefringence the complement of this model (in view of CPT status of LV terms), the Chern–Simons one, has been studied extensively in the literature and that $e\rightarrow e\gamma$ in vacuo were also explored in this context, both microscopically Kaufhold:2007qd and macroscopically Lehnert:2004hq . where the model is restricted to modified Maxwell theory. A systematic analysis incorporating dynamical aspects of LV for dimension-4 and -6 terms was carried out soon after in ref. Rubtsov:2012kb and it is the procedure for evaluation of rates described therein that we shall follow in our study. Of course, there also exist attempts similar to us, i.e., to consider radiation rates for the general case which converges quicker and simpler to phenomenology via (1) and is not (necessarily) bounded to a particular model; see for instance Martinez-Huerta:2017ulw , where the treatment is more general and, thus, technically more involved. However therein a priori the effect of LV in fermions is ignored; this is not case here along the next subsection.
@@ 2.2 Decay width for anomalous Cherenkov process of charges and the rate
where the residual rotation invariance implicit in eq. (1) has been used. [^7] We remind the reader, that this simplifying choice is used implicitly in most phenomenological analyses of LV, say, in He:2022jdl ; Li:2022ugz , as is also the cases of some noneffective field-theoretical QG models including the ones based on brane theories discussed below; otherwise spatial anisotropies enter (5) via direction-$i$-dependent $\eta_{i},\xi_{i}$ (for implications of broken O(3) symmetry see, e.g., Mattingly:2005re ). For the energy well above the threshold of interest here, the lepton masses can be safely neglected so that the electron/positron dispersion further simplifies to $E^{2}\simeq{\bf p}^{2}(1+\delta+\eta p/M)$ and, for the outgoing particle, $E^{\prime 2}\simeq{\bf p}^{\prime 2}(1+\delta+\eta p^{\prime}/M)$.
@@
$J=\int\frac{\textrm{d}k_{p}\textrm{d}p_{\perp}^{\prime 2}}{pk_{p}(1-k_{p})}2\delta\Bigl{(}\varpi(k_{p})-\frac{p_{\perp}^{\prime 2}}{pk_{p}(1-k_{p})}\Bigr{)}[EE^{\prime}-({\bf p}\cdot{\bf p}^{\prime})],$ (13)
@@
$\varpi(k_{p})=\delta\lvert{\bf p}\rvert k_{p}-\frac{\xi}{M}{\bf p}^{2}k_{p}^{2}+\frac{\eta}{M}{\bf p}^{2}(2k_{p}-k_{p}^{2}).$ (14)
@@
$J=2\int\textrm{d}p_{\perp}^{\prime 2}\int$ $\,\textrm{d}k_{p}\frac{1}{k_{p}(1-k_{p})}\delta\Bigl{(}p\varpi-\frac{p_{\perp}^{\prime 2}}{k_{p}(1-k_{p})}\Bigr{)}$
×$\times\Bigl{[}\delta p^{2}(1-k_{p})+\frac{\eta}{2M}p^{3}(1-k_{p})(2-k_{p})+\frac{p_{\perp}^{\prime 2}}{2(1-k_{p})}\Bigr{]},$ (15)
which despite its apparently complicated form, is just an elementary integration. Performing the integral over $p_{\perp}^{\prime}$, we obtain the differential emission rate:
@@
One observes that the rate is suppressed by the parameters describing LV. Notice however, that the suppression does not come from the phase space integration, which reduces to the integral over $k_{p}$, as below. Instead, it is due to the smallness of the matrix element for the nearly collinear kinematics realized in the decay.
@@
$\varpi(k_{p})\geq 0.$ (17)
This implies that, while the $k_{p}$-range is of the standard form (7), for some $k_{p}^{\max}$ specified by this condition, it will actually depend on the hierarchies among the LV parameters. For our purpose of showing that the decay is so fast that in particular in the above-mentioned cases of our interest (i.e., either $\xi$ or $\eta$ is zero) it is practically governed by pure kinematics classifying all possibilities is not relevant for us. In fact, it suffices to consider the simplest configuration $0\leq\delta$, $0\leq\eta$, $\xi\leq\eta$, which is appropriate for achieving our aim. In this case the integral is over the whole range $0\leq k_{p}\leq 1$, yielding,
@@
This result is exact with the only approximation being $m_{e}=0$. For one of the most useful characteristic of the process, the rate of the energy decrease by the electron, c.f. (11), one multiplies the integrand of $J$ by $-Ek_{p}$ and obtains,
@@
For $\xi=0$ and $0<\eta$ (or $\eta=0$, $0>\xi$) of its generic order ${\cal O}(1)$ with $\delta\approx 0$, $\tau\sim 8\alpha_{e}^{-1}M/p^{2}$ and we can see that the difference in the effective radiation time $\tau$ between these two cases is only up to an ${\cal O}(1)$ numerical factor.
Consider ultrahigh-energy electrons, for which one takes the highest energy (indirectly) detected by LHAASO detectors LHAASO:2021cbz , $E\sim 2$ PeV. As mentioned, they are widely regarded as the origin of the observed UHE IC photons of the Crab Nebula. As far as the fine structure constant is concerned we use its value at low energy, 1/137, neglecting its running coupling nature as it does not affect our result significantly. From eq. (20),
@@
where 10 TeV is the order of the threshold for ${\cal O}(1)$ $\eta$ (or $\xi$) as required by eqs. (2), (3). Then for $p={\cal O}(\textrm{PeV})$, $\tau$ is just $\sim 10^{-15}$ seconds, implying that the initial electron/positron loses a significant fraction of its excess energy really quickly, corresponding to propagation over a distance scale of micrometers. Since it certainly travel distances far larger than that, we must demand that the threshold for the vacuum Cherenkov effect is above the maximal energy of any charged lepton known to be stable, regardless of whether the emitted photon is soft or hard. This is precisely the condition utilized for setting bounds upon LV via such effects, thus constraining $\xi$ or $\eta$ to at least of ${\cal O}(10^{-6})$ He:2022jdl ; Li:2022ugz . If one adopts such improved constraints, the decay rate will be $\sim 10^{6}$ times smaller and particles radiate ‘all’ its energy down to thresholds in a slightly longer time of about 100 picoseconds, which are still short enough, that on any relevant astrophysical timescales electrons of PeV energies will almost completely decay any time it takes place.
We therefore conclude that for a general class of Lorentz-violating dispersion relations, of the type (1), in all cases of interest the ensuing electron Cherenkov process in vacuo is an efficient reaction, once kinematically allowed, as justified by the rapidity with which the decay rate becomes large above thresholds. It may be considered for all practical purposes, as happening instantaneously and the loss rate $\dot{E}\sim e^{2}E^{3}/M$ is high enough that one does not make any relevant mistakes if just exploiting the absence of thresholds for the purpose of limiting certain types of scales of the violation. Hence in particular, the stringent bound imposed by simply looking at the threshold condition (of the hard Cherenkov effect) on the linear LV scale entailed by a modified subluminal photon dispersion, $E_{\textrm{LV}\gamma}=M/(-\xi)^{-1}\gtrsim 1.2\times 10^{25}$ GeV (if $\eta=0$He:2022jdl is indeed reliable at least as long as we work in the realm of the purely phenomenological LV extension to the canonical relativistic dispersions. But, as we shall see in the next section (section 3) there exist QG scenarios where such a constraint do not apply and where there would be compatibility Li:2021gah with the hint of light-speed variation from time delay measurements of GRBs Shao:2009bv that require a value of order $10^{17}$ GeV for the same scale parameter.

### v2

@@ Abstract
Keywords:
D-Branes, String Models, Violation of Lorentz and/or CPT Symmetry
arxiv: 2505.06121
@@ 1 Introduction and summary
The discovery of PeV photons reported by the Large High Altitude Air Shower Observatory (LHAASO) from PeVatrons LHAASO:2021gok ; Li:2021tcw ; LHAASO:2023uhj including the ancient Crab Nebula LHAASO:2021cbz , kicks off a new era in astronomy, providing a unique opportunity to probe the potential breakdown of Lorentz invariance (LI) Li:2025stf ; AlvesBatista:2023wqm ; He:2022gyk arising from quantum gravity (QG) analysis. High-energy violation of this symmetry (i.e., Lorentz violation (LV)) leads to particle-energy ($E$) dependent physics, that is suppressed by the gravitational scale $M\sim M_{P}\approx 10^{19}$ GeV where $M_{P}$ is the Planck mass. As a result, for extreme energies reached in astrophysical processes, its consequences could be identified in the current observations. The question arises, then, what information we can get from LHAASO experiments of ultrahigh-energy (UHE) $\gamma$-rays and particularly, from $\sim 1.1$ PeV Crab photon. Recent studies He:2022jdl , carrying further a previous analysis Li:2022ugz by the present authors, pointed out that the (indirect) observations of UHE electrons assumed to produce the PeV emission of the Crab severely limit minimally suppressed LV dispersion corrections, i.e., of ${\cal O}(E/M)$, to both electrons and photons.
Such bounds were placed via considerations of the vacuum Cherenkov process, $e\rightarrow e\gamma$, which is forbidden in LI quantum electrodynamics (QED) since the speed of a fast moving charge is always less than that of photon propagation. With LV it can be Coleman:1997xq an important energy-loss channel for the most high-energetic particle. As often adopted in the literature to parameterize LV (see Li:2025stf ), also in the same spirit as He:2022jdl ; Li:2022ugz , a priori most phenomenological studies converge in assuming a generic, nonstandard dispersion relation:
@@
where $(E_{I},{\bf p}_{I})$ is the four-momentum associated with a given particle $I$ of (rest) mass $m_{I}$, and the modulus of the spatial momentum is sometimes denoted, below, by $p\coloneqq\lvert{\bf p}\rvert$. [N.B.: the notation follows prior studies of He:2022jdl ; Li:2022ugz .] $\ell_{I,n}\coloneqq s_{I,n}/(E_{\textrm{LV}I,n})^{n}\equiv-\xi_{I,n}/M^{n}$ defines the scale of the $n$th-order LV physics beneath, [^3] Here, the dimensionless LV parameters $\xi_{I,n}$ are defined to simplify some of the expressions below, albeit they may be absorbed into redefinition of the mass $M$, and, then give the (positive) LV energy scale, $E_{\textrm{LV}I,n}$ by keeping only the LV sign factor $s_{I,n}(=+1$ or 1)-1) explicitly. while $\delta^{(I\,=\,e)}$ parameterizes the strength of ${\cal O}(p^{2})$ electron LV that has been constrained to be extremely small, $\lvert\delta\rvert<10^{-20}$ Li:2022ugz , implying that for astrophysically relevant processes, ${\cal O}(1/M^{n})$ terms dominate.
For example, in a linear LV model ($n=1$, in which case this suffix $n$ is omitted) often considered, modified dispersion (1) induces an energy-dependent velocity of propagation in vacuo for both the electromagnetic waves, $\Delta v_{\gamma}\coloneqq\partial E_{\gamma}/\partial p_{\gamma}-1\simeq-\,(\ell E)_{\gamma}$ and electron waves, $\Delta v_{e}\simeq-[m_{e}^{2}/(2E_{e}^{2})+(\ell E)_{e}]$, assuming vanishing $\delta$, with $s_{\gamma/e}=+1$ (1-1) for subluminal (superluminal) propagation relative to low-energy photons that are not supposed to be affected significantly by the LV effects, and that will travel at the conventional speed of light $c$ (already set to be unity). [^4] Such nontrivial propagation properties can be found from concrete models. For example within effective field theory approach to modified QED, models with operators of dimension-5 are compatible with the case of exhibiting birefringence (i.e., having Myers:2003fd helicity-dependent $\ell_{\gamma/e}$) while (Liouville Ellis:1992eh ) string inspired foamy space-time ‘medium’ Mavromatos:2002re based on D-brane recoil Ellis:1999uh dictates a subluminal refraction for photons (and not electrons, c.f. $\ell_{e}=0$ at tree-level), see section 3. Aside from such propagation effects in vacuo there is the possibility of peculiar novel (threshold) reactions beyond the standard model, like $e\rightarrow e\gamma$, provided the energy–momentum conservation relation is kept intact, an assumption usually adopted in a purely phenomenological formulation of LV, e.g., (1), owing to a lack of any particular prior knowledge.
@@
Quite interestingly, the absence of hard Cherenkov threshold severely constrains subluminal photon dispersions with linear LV modification, an effect compatible with a previous hint of light-speed variation in gamma-ray burst (GRB) photon flight time delays Shao:2009bv , that require $E_{\textrm{LV}\gamma}\sim 10^{17}$ GeV, leading to a saturation of the latter. This motivates us to reconsider the bound from (3). In fact, only if we confirm that electrons lose energy extremely quickly through this process can we reliably apply the LV bounds derived from the threshold considerations above. To date, this issue has been examined only for specific models of LV Altschul:2007tn ; Klinkhamer:2008ky ; Anselmi:2011ae ; Rubtsov:2012kb , while the vacuum Cherenkov rate within purely phenomenological approaches, such as those described by (1), remains unclear. In the first part of this paper, we aim to briefly discuss this topic. Above the threshold, as we will demonstrate, the reaction rate is significant, allowing straightforward kinematic threshold analyses, as seen in refs. He:2022jdl ; Li:2022ugz , to yield reliable constraints.
However, at the kinematic level, permitting electron Cherenkov effect in vacuo requires two conditions which may need re-evaluation in certain QG frameworks: i) the adoption of LV-modified dispersion relations in eq. (1), and ii) the preservation of energy–momentum conservation. If either condition fails, ‘electron decay’ through such an anomalous channel could be ruled out (much like in the standard model, though for entirely different reasons), despite the presence of LV in the electron and/or photon sectors. This scenario may actually describe certain (solitonic) string/brane models of space-time foam, which we will examine more closely later in this work.
@@ 2.1 A note on the calculation of rates in Lorentz violation contexts
[…]
$\overline{\sum}\lvert{\cal M}^{\textrm{LV}}(\{p\})\rvert^{2}=\overline{\lvert{\cal M}(\{p\})\rvert^{2}}\,\bigl[1+{\cal O}(E/M)\bigr],$ (4)
@@
Additionally, this section focuses solely on illustrating the calculation of the Cherenkov rate to support the estimation of bounds derived from kinematic thresholds. This complements recent studies in He:2022jdl ; Li:2022ugz , and for consistency, we adopt the same theoretical framework as those studies. Consequently, we neglect potential effects arising from (4), assuming LI matrix element. Nonetheless as mentioned above, such reaction rates have been calculated in many earlier works, to which we refer the reader for details, and where one considers in general a concrete theory so that the dynamical effects (c.f. (4)) can be treated properly. For example, modifications of phase-space integrals and to the functional form of external states were studied in Jacobson:2005bg in the context of QED with dimension-5 field operators of which the dispersion induced is of $n=1$ type Myers:2003fd thus bears a similarity with the one considered here. Later, spin sums over external states for models with operators of higher dimensions were noted in Anselmi:2011ae , which built on earlier calculations of Klinkhamer:2008ky  [^6] Albeit severely constrained via the absence of detectable birefringence the complement of this model (in view of CPT status of LV terms), the Chern–Simons one, has been studied extensively in the literature and $e\rightarrow e\gamma$ in vacuo were also explored in this context, both microscopically Kaufhold:2007qd and macroscopically Lehnert:2004hq . where the model is restricted to modified Maxwell theory. A systematic analysis incorporating dynamical aspects of LV for dimension-4 and -6 terms was carried out soon after in ref. Rubtsov:2012kb and it is the procedure for evaluation of rates described therein that we shall follow in our study. Of course, there also exist attempts similar to us, i.e., to consider radiation rates for the general case which converges quicker and simpler to phenomenology via (1) and is not (necessarily) bounded to a particular model; see for instance Martinez-Huerta:2017ulw , where the treatment is more general and, thus, technically more involved. However therein a priori the effect of LV in fermions is ignored; this is not the case here along the next subsection.
@@ 2.2 Decay width for anomalous Cherenkov process of charges and the rate
where the residual rotation invariance implicit in eq. (1) has been used. [^7] We remind the reader, that this simplifying choice is used implicitly in most phenomenological analyses of LV, say, in He:2022jdl ; Li:2022ugz , as is also the cases of some noneffective field-theoretical QG models including the ones based on brane theories discussed below; otherwise spatial anisotropies enter (5) via direction-$i$-dependent $\eta_{i},\xi_{i}$ (for implications of broken O(3) symmetry see, e.g., Mattingly:2005re ). For the energy well above the threshold of interest here, the lepton masses can be safely neglected so that the electron/positron dispersion further simplifies to $E^{2}\simeq{\bf p}^{2}(1+\delta+\eta p/M)$, and for the outgoing particle $E^{\prime 2}\simeq{\bf p}^{\prime 2}(1+\delta+\eta p^{\prime}/M)$. We emphasize that there is no helicity or spin dependence of the LV terms in our oversimplified formalism (5), i.e., $\eta$ and $\xi$ are assumed to be either positive or negative (or equivalently, $s_{e/\gamma}$ takes only one value between 1-1 and +1+1). Otherwise, for example, there would be photon birefringence which has been severely constrained by cosmological measurements Li:2025stf ; AlvesBatista:2023wqm ; He:2022gyk .
@@
$J=\int\frac{\textrm{d}k_{p}\textrm{d}p_{\perp}^{\prime 2}}{pk_{p}(1-k_{p})}2\delta\Bigl(F(k_{p})-\frac{p_{\perp}^{\prime 2}}{pk_{p}(1-k_{p})}\Bigr)[EE^{\prime}-({\bf p}\cdot{\bf p}^{\prime})],$ (13)
@@
$F(k_{p})=\delta\lvert{\bf p}\rvert k_{p}-\frac{\xi}{M}{\bf p}^{2}k_{p}^{2}+\frac{\eta}{M}{\bf p}^{2}(2k_{p}-k_{p}^{2}).$ (14)
@@
$J=2\int\textrm{d}p_{\perp}^{\prime 2}\int$ $\,\textrm{d}k_{p}\frac{1}{k_{p}(1-k_{p})}\delta\Bigl(pF(k_{p})-\frac{p_{\perp}^{\prime 2}}{k_{p}(1-k_{p})}\Bigr)$
$\times\Bigl[\delta p^{2}(1-k_{p})+\frac{\eta}{2M}p^{3}(1-k_{p})(2-k_{p})+\frac{p_{\perp}^{\prime 2}}{2(1-k_{p})}\Bigr],$ (15)
which despite its apparently complicated form, is just an elementary integration. Performing the integral over $p_{\perp}^{\prime}$, then, we obtain the differential emission rate:
@@
One observes that the rate is suppressed by the parameters describing LV. This suppression stems from the phase space integration, which is identically zero in the LI case and remains small as long as LV is small, because the available phase space is very limited for the nearly collinear kinematics realized in the decay.
@@
$F(k_{p})\geq 0.$ (17)
This implies that, while the $k_{p}$-range is of the standard form (7), for some $k_{p}^{\max}$ specified by this condition, it will actually depend on the hierarchies among the LV parameters. For our purpose of showing that the decay is so fast that in particular in the above-mentioned cases of our interest (i.e., either $\xi$ or $\eta$ is zero) it is practically governed by pure kinematics, classifying all possibilities is not relevant for us. In fact, it suffices to consider the simplest configuration $0\leq\delta$, $0\leq\eta$, $\xi\leq\eta$, which is appropriate for achieving our aim. In this case the integral is over the whole range $0\leq k_{p}\leq 1$, yielding,
@@
This result is exact with the only approximation being $m_{e}=0$. For one of the most useful characteristic of the process, the rate of the energy decrease by the electron, c.f. (11), one multiplies the integrand of $J$ by $Ek_{p}$ and obtains,
@@
For $\xi=0$ and $0<\eta$ (or $\eta=0$, $0>\xi$) of its generic order ${\cal O}(1)$ with $\delta\approx 0$, $\tau\sim 8\alpha_{e}^{-1}M/p^{2}$ and we can see that the difference in the effective radiation time $\tau$ between these two cases is only up to an ${\cal O}(1)$ numerical factor. While our primary focus is on the linearly Planck-scale suppressed modifications to the dispersion relations (of the type arising from certain string inspired scenarios discussed in section 3), we note that qualitatively similar features have been observed for QED with dimension-4 LV modifications Altschul:2007tn ; Klinkhamer:2008ky : as demonstrated through detailed analytical calculations in ref. Altschul:2007tn , the energy loss rates from the soft and hard parts of the Cherenkov photon spectrum are comparable in magnitude in this theory. It should be stressed that the ${\cal O}(E/M)$ dispersion modifications correspond to dimension-5 terms Myers:2003fd within the framework of effective field theories, though our setting (5) further assumes helicity/spin independence.
Consider ultrahigh-energy electrons, for which one takes the highest energy (indirectly) detected by LHAASO detectors LHAASO:2021cbz , $E\sim 2$ PeV. As mentioned, they are widely regarded as the origin of the observed UHE IC photons of the Crab Nebula. From eq. (20), therefore we have,
@@
where 10 TeV is the order of the threshold for ${\cal O}(1)$ $\eta$ (or $\xi$) as required by eqs. (2), (3). Then for $p={\cal O}(\textrm{PeV})$, $\tau$ is just $\sim 10^{-15}$ seconds, implying that the initial electron/positron loses a significant fraction of its excess energy really quickly, corresponding to propagation over a distance scale of micrometers. Since it certainly travels distances far larger than that, we must demand that the threshold for the vacuum Cherenkov effect is above the maximal energy of any charged lepton known to be stable, regardless of whether the emitted photon is soft or hard. This is precisely the condition utilized for setting bounds upon LV via such effects, thus constraining $\xi$ or $\eta$ to at least of ${\cal O}(10^{-6})$ He:2022jdl ; Li:2022ugz . If one adopts such improved constraints, the decay rate will be $10^{6}$ times smaller and the particle radiates ‘all’ its energy down to thresholds in a slightly longer time of about 100 picoseconds, which are still short enough, that on any relevant astrophysical timescales electrons of PeV energies will almost completely decay any time it takes place.
We therefore conclude that for a general class of Lorentz-violating dispersion relations, of the type (1), in all cases of interest the ensuing electron Cherenkov process in vacuo is an efficient reaction, once kinematically allowed, as justified by the rapidity with which the decay rate becomes large above thresholds. It may be considered for all practical purposes, as happening instantaneously and the loss rate $\dot{E}\sim e^{2}E^{3}/M$ is high enough that one does not make any relevant mistakes if just exploiting the absence of thresholds for the purpose of limiting certain types of scales of the violation. Hence in particular, the stringent bound imposed by simply looking at the threshold condition (of the hard Cherenkov effect) on the linear LV scale entailed by a modified subluminal photon dispersion, $E_{\textrm{LV}\gamma}=M/(-\xi)^{-1}\gtrsim 1.2\times 10^{25}$ GeV (if $\eta=0$He:2022jdl is indeed reliable at least as long as we work in the realm of the purely phenomenological LV extension to the canonical relativistic dispersions. But, as we shall see in the next section (section 3) there exist QG models where such a constraint does not apply and where there would be compatibility Li:2021gah with the hint of light-speed variation from time delay measurements of GRBs Shao:2009bv that require a value of order $10^{17}$ GeV for the same scale parameter.